Session: Jun 2014 (session 2)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Noto (m.paterno at Thu Jun 19 00:57:22 2014) Success:

Metsahovi (M. Uunila at Thu Jun 19 04:39:58 2014) Success:

Yebes (Angel V. Lozano at Thu Jun 19 05:46:51 2014) Success:
Change Disk on Scan Noo100

Svetloe (Slobojaninova N. at Thu Jun 19 10:07:43 2014) Success:

Zelenchukskaya (Kononenko V. at Thu Jun 19 10:08:35 2014) Success:
Scans from No0136 to No0226 missed - pause for intensive geo session

Badary (Lysakova, Naumov at Thu Jun 19 10:09:47 2014) Success:
Pause for intensive geo session from 18:30 to 21:28 (scans no0133-no0251).

Jodrell_Bank (Jimmy Cullen at Thu Jun 19 11:37:14 2014) Failure:RECE
Unknown problem, but we suspect a fault in the receiver

Torun (K. Borkowski at Fri Jun 20 10:20:28 2014) Success:
No known problems.

Onsala (Jun Yang at Mon Jun 23 07:04:16 2014) Success:
No known technical problems. The Tsys data were quit high. Note that the atmospheric opacity was already corrected in the log and antabfs files as Chopper wheel calibration was used. The average opacity was 0.49 during the experiment.

Effelsberg (Uwe Bach at Mon Jun 23 14:04:01 2014) Success: